OpenStudy (anonymous):

what is a2+b2-c2=abcos(theta)?

OpenStudy (amistre64):

missing a 2 i think; looks very law of cosine ish

OpenStudy (ash2326):

\(a^2+b^2-c^2=2ab \cos \theta\)